Solve this equation:
Log3(x+2) - log3x=2
(The 3's are suppose to be lower than I typed them )

Answer by jim_thompson5910(35256)

I'm not going to type the '3's, but keep in mind that each log that is typed is a log base 3.

log(x+2) - log(x) = 2

log[(x+2)/(x)] = 2

(x+2)/(x) = 3^2

(x+2)/(x) = 9

x+2 = 9x

2 = 9x-x

2 = 8x

2/8 = x

1/4 = x

x = 1/4

So the final answer is x = 1/4